OpenStudy (anonymous):

6m^3n^4 divided by 28mn^5

OpenStudy (unklerhaukus):

\[\frac{6m^3n^4}{28mn^5}=\frac{2\times3}{2^2\times7}m^{3-1}n^{4-5}\]

OpenStudy (anonymous):

answers are:A.\[3m^2\div 14n\]. B.\[m^2\div 22n\]. C.\[m^4n^9\div 4\]. D.\[-22m^2n\]

OpenStudy (unklerhaukus):

\[...=\frac{3m^2}{2\times7n}=...\]

OpenStudy (anonymous):

thanks